OpenStudy (anonymous):

check my work please. estimate using linearization 1/0.99

OpenStudy (anonymous):

i have f(0.99)=1/1-1/(1)^2 * (0.99-1) ans=1.01

jimthompson5910 (jim_thompson5910):

what is the full problem?

OpenStudy (anonymous):

it just says estimate the value of 1/0.99 using linearization

jimthompson5910 (jim_thompson5910):

oh f(x) = 1/x is your original function

OpenStudy (anonymous):

yep

jimthompson5910 (jim_thompson5910):

f ' (x) = -1/(x^2) ok

OpenStudy (anonymous):

my f'(x) is -1/x^2

jimthompson5910 (jim_thompson5910):

everything looks good because the linearization function L(x) is L(x) = f(a) + f ' (a)*(x - a) where a = 1 in this case (since f(1) = 1 is an easy number to compute) and I'm also getting L(0.99) = 1.01

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

i got a question though

OpenStudy (anonymous):

for example if you have 1/2.99

OpenStudy (anonymous):

is your a = 3?

jimthompson5910 (jim_thompson5910):

then the old L(x) is a lousy approximation which means you'd have to change to something like a = 3, yes correct

jimthompson5910 (jim_thompson5910):

1/x is easy to work with for integer values of x

OpenStudy (anonymous):

i see. thanks for your help :D

jimthompson5910 (jim_thompson5910):

np